# Compressed Image Generation with Denoising Diffusion Codebook Models

### v3

@@ 2 Related Work
Image compression has seen significant progress in recent years, with the penetration of neural networks to this domain. Neural methods range from constructing specialized architectures (Ballé et al., 2017; Zhu et al., 2022; Jiang et al., 2023; Jiang & Wang, 2023) to relying on different generative models such as GANs (Mentzer et al., 2020; Muckley et al., 2023; Iwai et al., 2024) and VAEs (Theis et al., 2017). Recent compression methods leverage DDMs and offer high perceptual quality results, by training models from scratch (Yang & Mandt, 2023; Ghouse et al., 2023), fine-tuning existing models (Careil et al., 2023; Körber et al., 2024), or using pre-trained DDMs in a zero-shot manner (without further training) (Theis et al., 2022; Elata et al., 2024). Current solutions in the latter category are highly computationally demanding, either due to their communication schemes (e.g., reverse channel coding (Theis & Ahmed, 2022; Theis et al., 2022)) or their need to perform thousands of denoising operations (Elata et al., 2024). Our work falls into this last category, offering a novel and highly effective compression scheme with a fast bit-stream communication method and computational demands that match standard use of DDMs.
@@
Recent works have explored discrete generative modeling of continuous data distributions. These employ various discrete representations, such as vector quantized latent tokens (Wu et al., 2024) or hierarchical modeling schemes that gradually refine each generated sample (Yang, 2024). DDCM offers a new such discrete generative framework, building on the exceptional achievements of DDMs.
@@
The task of compressed image generation (generating images directly in their compressed form) has been previously explored. Kang et al. (2019) trained an unconditional Generative Adversarial Network (GAN) (Goodfellow et al., 2014) to synthesize JPEG representations. Rajesh et al. (2023) trained a text-conditional GAN in the JPEG domain. In this work, we use DDMs instead of GANs and introduce a novel compressed image representation space different than that of JPEG. Our approach is compatible with any pre-trained DDM without requiring additional training.
@@ 4 Denoising Diffusion Codebook Models
Figure 3: Comparing DDPM with DDCM for different codebook sizes $K$. Interestingly, DDCM with $K=64$ achieves similar FID to DDPM, suggesting that the continuous representation space of DDPM (DDCM with $K=\infty$) is highly redundant. We use a class-conditional ImageNet model ($256\times 256$) for pixel space, and the text-conditional SD 2.1 model ($768\times 768$) for latent space, with prompts from MS-COCO. The $K$ axis is in log-scale.
@@
While DDPM is equivalent to DDCM with $K=\infty$, the first question we address is whether DDCM maintains the synthesis capabilities of DDPM for relatively small $K$ values. We compare the performance of DDPM with that of DDCM using $K\in\{2,4,8,16,64\}$ for sampling from pre-trained pixel and latent space models. We compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the generation performance. In App. A we report additional metrics and provide qualitative comparisons. For pixel space generation, we use a class-conditional DDM trained on ImageNet $256\times 256$ (Deng et al., 2009; Dhariwal & Nichol, 2021), and apply classifier guidance (CG) (Dhariwal & Nichol, 2021) with unit scale. We use the 50k validation set of ImageNet as the reference dataset, and sample 10k class labels randomly to generate the images. For latent space, we use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) trained on $768\times 768$ images and apply classifier-free guidance (CFG) with scale 3 (equivalent to $w=2$ in (Ho & Salimans, 2021)). As the reference dataset, we randomly select 10k images from MS-COCO (Lin et al., 2014; Chen et al., 2015) along with one caption per image, and use those captions as prompts for sampling.
@@ 5 Image Compression with DDCM
Figure 4: Qualitative image compression results. The presented images are taken from the Kodak24 ($512\times 512$) dataset. Our codec produces highly realistic outputs, while maintaining better fidelity to the original images compared to previous methods.
@@
The bit rate of this approach is determined by the size of the codebooks $K$, and the number of sampling timesteps $T$. Specifically, the bit-stream length is given by $(T-1)\log_{2}(K)$. Therefore, the bit rate can be reduced by simply decreasing the number of codebooks, or by using a smaller number of timesteps at generation, e.g., by skipping every other step, or by using specific timestep intervals (see Sec. B.3). In the approach described so far, the length of the bitstream increases logarithmically with $K$, making it computationally demanding to increase the bit rate. For instance, even for $K=8192$, $T=1000$ and $768\times 768$ images our BPP is approximately 0.022. Thus, to produce higher bit rates, we propose to refine the noise selected at timestep $i$ by employing matching pursuit (MP) (Mallat & Zhang, 1993). Specifically, at each step $i$, we construct the chosen noise as a convex combination of $M$ elements from ${\mathcal{C}}_{i}$, gathered in a greedy fashion to best correlate with the guiding residual ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$ (as in Eq. 7). The resulting convex combination involves $M-1$ quantized scalar coefficients, chosen from a finite set of $C$ values taken from $[0,1]$. Therefore, the resulting length of the bit-stream is given by $\smash{(T-1)(\log_{2}(K)M+C(M-1))}$, such that $M=1$ is similar to our standard compression scheme, and the length of the bit-stream increases linearly with $M$ and $C$. We apply this algorithm when the absolute bits number crosses $(T-1)\cdot\log_{2}(2^{13})$. Further details are available in Sec. B.4.
@@
We evaluate our compression method on Kodak24 (Franzen, 1999), DIV2K (Agustsson & Timofte, 2017), ImageNet 1K $256\times 256$ (Deng et al., 2009; Pan et al., 2020), and CLIC2020 (Toderici et al., 2020). For all datasets but ImageNet, we center crop and resize all images to $512\times 512$. We compare to numerous competing methods, both non-neural and neural, and both zero-shot, fine-tuning based, and training based approaches. For the ImageNet dataset, we use the unconditional pixel space ImageNet $256\times 256$ model of Dhariwal & Nichol (2021), and compare our results to BPG (Bellard, 2018), HiFiC (Mentzer et al., 2020), IPIC (Xu et al., 2024), and two PSC (Elata et al., 2024) configurations, distortion-oriented (PSC-D) and perception-oriented (PSC-P). For all other datasets, we use SD 2.1 $512\times 512$ (Rombach et al., 2022) and compare to BPG, PSC-D, PSC-P, ILLM (Muckley et al., 2023), PerCo (SD) (Körber et al., 2024; Careil et al., 2023), and twoCRDR (Iwai et al., 2024) configurations, distortion-oriented (CRDR-R) and perception-oriented (CRDR-R). PSC shares the same pre-trained model as ours, while PerCo (SD) requires additional fine-tuning. For our method, we apply SD 2.1 unconditionally, as we saw no improvement by adding prompts (see further details in Sec. B.5). We assess our method for several options of $T$, $K$, $M$, and $C$ to control the bit rate. See further details in App. B. We evaluate distortion with PSNR and LPIPS (Zhang et al., 2018) and perceptual quality with FID (Bińkowski et al., 2018). For ImageNet, FID is computed against the entire 50k $256\times 256$ validation set. For the smaller datasets we follow Mentzer et al. (2020) and compute the FID over extracted image patches. Specifically, for DIV2K and CLIC2020 we extract $128\times 128$ sized patches, and for Kodak we use $64\times 64$.
@@
As shown in Fig. 5, our compression scheme dominates previous methods on the rate-perception-distortion plane (Blau & Michaeli, 2019) for lower bit rates, surpassing both the perceptual quality (FID) and distortion (PSNR and LPIPS) of previous methods. For instance, our FID scores are lower than those of all other methods at around 0.1 BPP, while, for the same BPP, our distortion performance is better than the perceptually-oriented methods (e.g., PerCo, PSC-P, and IPIC). However, our method under-performs at the highest bit rates, especially when using SD. This is expected due to the performance ceiling entailed by the pre-trained encoder-decoder of SD (Körber et al., 2024; Elata et al., 2024). The qualitative comparisons in Fig. 4 further demonstrate our superior perceptual quality, where, even for extreme bit rates, our method maintains similarity to the original images in fine details. See App. B for more details and results.
@@
Figure 5: Compression quantitative evaluation. We compare the perceptual quality (FID) and distortion (PSNR, LPIPS) achieved at different BPPs. The image sizes of each dataset is denoted next to its name. Our method produces the best perceptual quality at most BPPs. Importantly, this is while we also attain lower distortion compared to perceptually-oriented methods (e.g., PSC-P and PerCo (SD)). For the three rightmost datasets, note that our approach, PSC-P, and PerCo (SD) use the latent space Stable Diffusion 2.1 model, so its VAE imposes a distortion bound. Thus, we report the distortion attained by simply passing the images through this VAE (dashed line).
@@ 6.1 Compressed Posterior Sampling for Image Restoration
Following (Chung et al., 2023; Wang et al., 2023b), we implement our method using the unconditional ImageNet $256\times 256$ DDM trained by Dhariwal & Nichol (2021). We fix $K=4096$ for all codebooks, resulting in a compressed bit-stream of approximately $0.183$ BPP for each generated image. We compare our method with DPS (Chung et al., 2023) and DDNM (Wang et al., 2023b) on two noiseless tasks: colorization and 4×4$\times$4 × super-resolution (using the bicubic kernel). We evaluate these methods using their official implementations and the same DDM. We additionally compress the outputs of DPS and DDNM to assess whether such a naive approach would yield better results. To do so, we adopt our proposed compression scheme (from Sec. 5), employing the same unconditional ImageNet DDM and using $K=4096$ noises per codebook.
@@ 6.2 Compressed Real-World Face Image Restoration
Real-world face image restoration is the practical task of restoring any degraded face image, without any knowledge of the corruption process it has gone through (Wang et al., 2021; Gu et al., 2022; Wang et al., 2022; Zhou et al., 2022; Wang et al., 2023c; Lin et al., 2024; Yue & Loy, 2024; Chen et al., 2024b; Ohayon et al., 2024). We propose a novel method capable of optimizing any no-reference image quality assessment (NR-IQA) measure at test time (e.g., NIQE (Mittal et al., 2013)), without relying on gradients.
@@
[…]
@@ 7 Discussion
While our empirical results are encouraging, our work does not explain theoretically why DDCM sampling and our simple index selection strategies work so effectively. We encourage future works to investigate the principles behind the success of our methods. Moreover, when operating in latent space, our codec’s performance is bounded by the underlying VAE, particularly at higher bit rates. Our results at higher bit rates could also be improved through better approaches than our current matching pursuit inspired solution. Additionally, our compression efficiency could be improved through entropy coding of the selected indices, potentially reducing bit rates without sacrificing quality. Lastly, all DDCM-based solutions can be improved further by optimizing the codebooks, e.g., through dictionary learning. To conclude, our work demonstrates promising results across numerous tasks while providing opportunities for both theoretical analysis and practical improvements.
@@ Appendix A Additional DDCM Evaluation
Figure 8: Comparing DDPM with DDCM for different codebook sizes $K$. As in Sec. 4, DDPM and DDCM with $K=64$ (sometimes even $K=16$) achieve similar generative performance, suggesting that the continuous representation space of DDPM (DDCM with $K=\infty$) is highly redundant. We use a class-conditional ImageNet model ($256\times 256$) for pixel space, and the text-conditional SD 2.1 model ($768\times 768$) for latent space. The $K$ axis is in log-scale.
Figure 9: Qualitative comparison of sample quality and diversity between DDCM and DDPM. We generate multiple samples for each prompt, using the $768\times 768$ SD 2.1 model.
Figure 10: Qualitative comparison of sample quality and diversity between DDCM and DDPM. We generate multiple samples for each class, using the conditional $256\times 256$ ImageNet model.

### v4

@@ 2 Related Work
Image compression has seen significant progress in recent years, with the penetration of neural networks to this domain. Neural methods range from constructing specialized architectures (Ballé et al., 2017; Zhu et al., 2022; Jiang et al., 2023; Jiang & Wang, 2023) to relying on different generative models such as Generative Adversarial Networks (GANs) (Mentzer et al., 2020; Muckley et al., 2023; Iwai et al., 2024) and Variational Autoencoders (VAEs) (Theis et al., 2017). Recent compression methods leverage DDMs and offer high perceptual quality results, by training models from scratch (Yang & Mandt, 2023; Ghouse et al., 2023), fine-tuning existing models (Careil et al., 2023; Körber et al., 2024), or using pre-trained DDMs in a zero-shot manner (without further training) (Theis et al., 2022; Elata et al., 2024). Current solutions in the latter category are highly computationally demanding, either due to their communication schemes (e.g., reverse channel coding (Theis & Ahmed, 2022; Theis et al., 2022)) or their need to perform thousands of denoising operations (Elata et al., 2024). Our work falls into this last category, offering a novel and highly effective compression scheme with a fast bit-stream communication method and computational demands that match standard use of DDMs.
@@
Recent works have explored discrete generative modeling of continuous data distributions. These employ various discrete representations, such as vector quantized latent tokens (Wu et al., 2024) or hierarchical modeling schemes that gradually refine each generated sample (Yang, 2025). DDCM offers a new such discrete generative framework, building on the exceptional achievements of DDMs.
@@
The task of compressed image generation (generating images directly in their compressed form) has been previously explored. Kang et al. (2019) trained an unconditional GAN (Goodfellow et al., 2014) to synthesize JPEG representations. Rajesh et al. (2023) trained a text-conditional GAN in the JPEG domain. In this work, we use DDMs instead of GANs and introduce a novel compressed image representation space different than that of JPEG. Our approach is compatible with any pre-trained DDM without requiring additional training.
Compressed Conditional Image Generation.
Our work also addresses the problem of compressed conditional generation (see Sec. 6), in which the compressed output is generated in accordance with a given input condition, such as a text prompt, a noisy image, or other forms of guidance. In this context, Liu et al. (2021) proposed a method based on optimal transport, subject to an informational bottleneck constraint. Specifically, their approach involves computing an optimal transport map from the distribution of input conditions to the distribution of target signals, while constraining the entropy of an intermediate latent random variable (i.e., the code). This method requires both the inputs and outputs to reside within the same metric space. In contrast, our approach naturally accommodates more complex types of input-output pairs, such as those involving different modalities (e.g., text and image).
@@ 4 Denoising Diffusion Codebook Models
Figure 3: Comparing DDPM with DDCM for different codebook sizes $K$. Interestingly, DDCM with $K=64$ achieves similar FID to DDPM, suggesting that the continuous representation space of DDPM (DDCM with $K=\infty$) is highly redundant. We use a class-conditional ImageNet model ($\times 256$ 256) for pixel space, and the text-conditional SD 2.1 model ($\times 768$ 768) for latent space, with prompts from MS-COCO. The $K$ axis is in log-scale.
@@
While DDPM is equivalent to DDCM with $K=\infty$, the first question we address is whether DDCM maintains the synthesis capabilities of DDPM for relatively small $K$ values. We compare the performance of DDPM with that of DDCM using $K\in\{2,4,8,16,64\}$ for sampling from pre-trained pixel and latent space models. We compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the generation performance. In App. A we report additional metrics and provide qualitative comparisons. For pixel space generation, we use a class-conditional DDM trained on ImageNet $\times 256$ 256 (Deng et al., 2009; Dhariwal & Nichol, 2021), and apply classifier guidance (CG) (Dhariwal & Nichol, 2021) with unit scale. We use the 50k validation set of ImageNet as the reference dataset, and sample 10k class labels randomly to generate the images. For latent space, we use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) trained on $\times 768$ 768 images and apply classifier-free guidance (CFG) with scale 3 (equivalent to $w=2$ in (Ho & Salimans, 2021)). As the reference dataset, we randomly select 10k images from MS-COCO (Lin et al., 2014; Chen et al., 2015) along with one caption per image, and use those captions as prompts for sampling.
@@ 5 Image Compression with DDCM
Figure 4: Qualitative image compression results. The presented images are taken from the Kodak24 ($\times 512$ 512) dataset. Our codec produces highly realistic outputs, while maintaining better fidelity to the original images compared to previous methods.
@@
The bit rate of this approach is determined by the size of the codebooks $K$, and the number of sampling timesteps $T$. Specifically, the bit-stream length is given by $(T-1)\log_{2}(K)$. Therefore, the bit rate can be reduced by simply decreasing the number of codebooks, or by using a smaller number of timesteps at generation, e.g., by skipping every other step, or by using specific timestep intervals (see Sec. B.4). In the approach described so far, the length of the bitstream increases logarithmically with $K$, making it computationally demanding to increase the bit rate. For instance, even for $K=8192$, $T=1000$ and $\times 768$ 768 images our BPP is approximately 0.022. Thus, to produce higher bit rates, we propose to refine the noise selected at timestep $i$ by employing matching pursuit (MP) (Mallat & Zhang, 1993). Specifically, at each step $i$, we construct the chosen noise as a convex combination of $M$ elements from ${\mathcal{C}}_{i}$, gathered in a greedy fashion to best correlate with the guiding residual ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$ (as in Eq. 7). The resulting convex combination involves $M-1$ quantized scalar coefficients, chosen from a finite set of $C$ values taken from [0,1][0,1][ 0 , 1 ]. Therefore, the resulting length of the bit-stream is given by $\smash{(T-1)(\log_{2}(K)M+C(M-1))}$, such that $M=1$ is similar to our standard compression scheme, and the length of the bit-stream increases linearly with $M$ and $C$. We apply this algorithm when the absolute bits number crosses $(T-1)\cdot\log_{2}(2^{13})$. Further details are available in Sec. B.5.
@@
We evaluate our compression method on Kodak24 (Franzen, 1999), DIV2K (Agustsson & Timofte, 2017), ImageNet 1K $\times 256$ 256 (Deng et al., 2009; Pan et al., 2020), and CLIC2020 (Toderici et al., 2020). For all datasets but ImageNet, we center crop and resize all images to $\times 512$ 512. We compare to numerous competing methods, both non-neural and neural, and both zero-shot, fine-tuning based, and training based approaches. For the ImageNet dataset, we use the unconditional pixel space ImageNet $\times 256$ 256 model of Dhariwal & Nichol (2021), and compare our results to BPG (Bellard, 2018), HiFiC (Mentzer et al., 2020), IPIC (Xu et al., 2024), and two PSC (Elata et al., 2024) configurations, distortion-oriented (PSC-D) and perception-oriented (PSC-P). For all other datasets, we use SD 2.1 $\times 512$ 512 (Rombach et al., 2022) and compare to BPG, PSC-D, PSC-P, ILLM (Muckley et al., 2023), PerCo (SD) (Körber et al., 2024; Careil et al., 2023), and twoCRDR (Iwai et al., 2024) configurations, distortion-oriented (CRDR-R) and perception-oriented (CRDR-R). PSC shares the same pre-trained model as ours, while PerCo (SD) requires additional fine-tuning. For our method, we apply SD 2.1 unconditionally, as we saw no improvement by adding prompts (see further details in Sec. B.6). We assess our method for several options of $T$, $K$, $M$, and $C$ to control the bit rate. See further details in App. B. We evaluate distortion with PSNR and LPIPS (Zhang et al., 2018) and perceptual quality with FID (Bińkowski et al., 2018). For ImageNet, FID is computed against the entire 50k $\times 256$ 256 validation set. For the smaller datasets we follow Mentzer et al. (2020) and compute the FID over extracted image patches. Specifically, for DIV2K and CLIC2020 we extract $\times 128$ 128 sized patches, and for Kodak we use $\times 64$ 64.
@@
As shown on the rate-distortion and rate-perception planes in Fig. 5, our compression scheme dominates previous methods on the rate-perception-distortion tradeoff (Blau & Michaeli, 2019) for lower bit rates, surpassing both the perceptual quality (FID) and distortion (PSNR and LPIPS) of previous methods. For instance, our FID scores are lower than those of all other methods at around 0.1 BPP, while, for the same BPP, our distortion performance is better than the perceptually-oriented methods (e.g., PerCo, PSC-P, and IPIC). However, our method under-performs at the highest bit rates, especially when using SD. When using a latent DDM such as SD, we aim to compress the latent encoding of a given image, rather than the image itself. Since encoding and decoding an image typically leads to a distorted reconstruction, the distortion of the compression method is bounded by that of the encoder-decoder. This leads to a performance ceiling for any latent-space-based compression method (Körber et al., 2024; Elata et al., 2024). We therefore include in Fig. 5 the “SD 2.1 Encoder-Decoder bound”, which corresponds to the distortion resulting from encoding and decoding the original images using the VAE of SD 2.1, without any additional compression. The qualitative comparisons in Fig. 4 further demonstrate our superior perceptual quality, where, even for extreme bit rates, our method maintains similarity to the original images in fine details. See App. B for more details and results.
@@
Figure 5: Compression quantitative evaluation. We compare the perceptual quality (FID) and distortion (PSNR, LPIPS) achieved at different BPPs. The image sizes of each dataset is denoted next to its name. Our method produces the best perceptual quality at most BPPs. Importantly, this is while we also attain lower distortion compared to perceptually-oriented methods (e.g., PSC-P and PerCo (SD)). For the three rightmost datasets, note that our approach, PSC-P, and PerCo (SD) use the latent space Stable Diffusion 2.1 model, so its encoder-decoder imposes a distortion bound. Thus, we report the distortion attained by simply passing the images through this encoder-decoder (dashed line).
@@ 6.1 Compressed Posterior Sampling for Image Restoration
Following (Chung et al., 2023; Wang et al., 2023b), we implement our method using the unconditional ImageNet $\times 256$ 256 DDM trained by Dhariwal & Nichol (2021). We fix $K=4096$ for all codebooks, resulting in a compressed bit-stream of approximately 0.1830.1830.183 BPP for each generated image. We compare our method with DPS (Chung et al., 2023) and DDNM (Wang et al., 2023b) on two noiseless tasks: colorization and 4×4$\times$4 × super-resolution (using the bicubic kernel). We evaluate these methods using their official implementations and the same DDM. We additionally compress the outputs of DPS and DDNM to assess whether such a naive approach would yield better results. To do so, we adopt our proposed compression scheme (from Sec. 5), employing the same unconditional ImageNet DDM and using $K=4096$ noises per codebook.
@@ 6.2 Compressed Real-World Face Image Restoration
Real-world face image restoration is the practical task of restoring any degraded face image, without any knowledge of the corruption process it has gone through (Wang et al., 2021; Gu et al., 2022; Wang et al., 2022; Zhou et al., 2022; Wang et al., 2023c; Lin et al., 2024; Yue & Loy, 2024; Chen et al., 2024b; Ohayon et al., 2025). We propose a novel method capable of optimizing any no-reference image quality assessment (NR-IQA) measure at test time (e.g., NIQE (Mittal et al., 2013)), without relying on gradients.
@@
Note that this index selection rule is similar to that of our standard compression, replacing $\smash{{\mathbf{x}}_{0}}$ in Eq. 7 with $\smash{{\bm{r}}({\mathbf{y}})}$. This choice of indices in DDCM would lead to a reconstructed estimate of the MMSE solution ${\bm{y}}({\mathbf{y}})$, yielding blurry results with low distortion (Blau & Michaeli, 2018). In contrast, randomly picking a sequence of indices in DDCM would produce a high quality sample from the data distribution $p_{0}$. Therefore, we randomly choose $\smash{k_{i,P}\sim\text{Unif}(\{1,\ldots,K\})}$. Then, we use the DDM and compute $\smash{\hat{{\mathbf{x}}}_{0|i-1}}$ for each index $\smash{k\in\{k_{i,D},k_{i,P}\}}$ separately, denoting each result accordingly by $\hat{{\mathbf{x}}}_{0|i-1}^{(k)}$. The final index is picked to optimize the perception-distortion tradeoff via
@@ 7 Discussion
While our empirical results are encouraging, our work does not explain theoretically why DDCM sampling and our simple index selection strategies work so effectively. We encourage future works to investigate the principles behind the success of our methods. Moreover, when operating in latent space, our codec’s performance is bounded by the underlying VAE, particularly at higher bit rates. Our results at higher bit rates could also be improved through better approaches than our current matching pursuit inspired solution. Additionally, our compression efficiency could be improved through entropy coding of the selected indices, potentially reducing bit rates without sacrificing quality. Lastly, all DDCM-based solutions can be improved further by optimizing the codebooks, e.g., through dictionary learning.
While DDCM produces compressed image representations in the form of indices, it is worth noting that the codebook entries can be interpreted as discrete image tokens, which are chosen autoregressively. This implies that DDCM can be interpreted as an image tokenization method, forming a “language” for images. Such a perspective opens the door to intriguing possibilities, such as using the tokenized image representation as a natural text-conditioning mechanism for unconditional diffusion models. For instance, one could train a transformer to predict the sequence of codebook indices corresponding to an image given its text description. Such a transformer can then be used to generate text-conditional images, by predicting the corresponding sequence of DDCM codebook indices.
To conclude, our work demonstrates promising results across numerous tasks while providing opportunities for both theoretical analysis and practical improvements.
@@ Appendix A Additional DDCM Evaluation
Figure 8: Comparing DDPM with DDCM for different codebook sizes $K$. As in Sec. 4, DDPM and DDCM with $K=64$ (sometimes even $K=16$) achieve similar generative performance, suggesting that the continuous representation space of DDPM (DDCM with $K=\infty$) is highly redundant. We use a class-conditional ImageNet model ($\times 256$ 256) for pixel space, and the text-conditional SD 2.1 model ($\times 768$ 768) for latent space. The $K$ axis is in log-scale.
Figure 9: Qualitative comparison of sample quality and diversity between DDCM and DDPM. We generate multiple samples for each prompt, using the $\times 768$ 768 SD 2.1 model.
Figure 10: Qualitative comparison of sample quality and diversity between DDCM and DDPM. We generate multiple samples for each class, using the conditional $\times 256$ 256 ImageNet model.